OpenStudy (anonymous):

I don't understand how to do this problem? The volume of a sphere is 5,000 m3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (anonymous):

what's the graph

OpenStudy (anonymous):

i only have answer choices that are, 3,033 m2 62,832 m2 195 m2 1,517 m2

OpenStudy (anonymous):

we know that v=4/3pi*r^3=5000 you can from that equation to get r the you can use s=4poi*r^2 to get surface area

OpenStudy (anonymous):

s=4*pi*r^2